I need help with story problems
Solve the problem and show solution used......In a town election the winning candidate had 220 more votes than the loser. If 810 votes were cast in all how many did each candidate receive?
thanks
heather

Found 2 solutions by checkley71, ptaylor:
Answer by checkley71(8403)

X+(X+220)=810
X+X+220=810
2X=810-220
2X=590
X=590/2
X=295 VOTES FOR THE LOOSER.
X+220=515 VOTES FOR THE WINNER.
PROOF
295+515=810
810=810

Answer by ptaylor(2048)

Let x=number of votes for the loser
Then x+220=number of votes for the winner
Now we are told that:
loser votes + winner votes=810. So our equation to solve is:
x+x+220=810 subtract 220 from both sides
x+x+220-220=810-220 collect like terms
2x=590 divide both sides by 2
x=295---------------------------------number of votes for loser
x+220=295+220=515----------------------number of votes for winner
CK
295+515=810
810=810
and
295+220=515
515=515

Hope this helps---ptaylor